OpenStudy (anonymous):

What is the sum 1/g+2 +3/g+1 a. 3/g+3 b. g+3/(g+1)(g+2) c. 4g+7/(g+1)(g+2) d. 2g+3/(g+1)(g+2)

OpenStudy (anonymous):

g+2 and g+1 are the denominators of the fractions. To add the fractions together, you need to create a common denominator. In this case, how would you make that common denominator?

OpenStudy (anonymous):

1/g+2 +3/g+1 denominators are (g+2)and (g+1). LCD is (g+2)(g+1) 1/g+2 +3/g+1 =(g+1)/((g+1)(g+2)) + 3(g+2)/((g+1)(g+2)) =(g+1+3g+6)/((g+1)(g+2)) =the last step is left for you